# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ 3.3 Progress Allocation
In Algorithm 2, model TimePredict is used to estimates the time step (trajectory length) needed to transition from the current state $\mathbf{x}$ to the new state $\mathbf{x}^{\prime}$. This model is trained on the same trajectory data, that will be used to train the Diffusion model. It assumes that the trajectory length between two states follows a Gaussian distribution. A simple multilayer perceptron (MLP) is used as the backbone of TimePredict model, and it is trained to predict the mean and variance of the trajectory length using a negative log-likelihood loss function. To account for tasks with avoidance requirements, which may require longer trajectories, a scaling factor $\gamma$ is applied to the predicted mean trajectory length. This factor allows control over the algorithm’s conservativeness by adjusting the predicted trajectory length as needed.
@@ 3.4 Trajectory Generation
In the above part, a sequence of waypoints $\bar{\mathbf{s}}$ is obtained and the time intervals for each invariance progress have also been determined. The remaining task is to generate executable trajectories that connect the waypoints while ensuring the trajectories satisfy the invariance progresses. Specifically, the trajectory $\tau$ between two adjacent waypoints $(\mathbf{x}_{i},t_{i})$ and $(\mathbf{x}_{i+1},t_{i+1})$ must be feasible under the current system dynamics and satisfy the following conditions:
  1. 1.
    The trajectory length is $t_{i+1}-t_{i}+1$.
  2. 2.
    It starts at state $\mathbf{x}_{i}$ and ends at state $\mathbf{x}_{i+1}$.
  3. 3.
    It satisfies all invariance progresses active within the time interval $[t_{i},t_{i+1}]$.
We employ the diffusion model to solve this conditional trajectory generation problem since it can learn the distribution of system trajectories from system trajectory data, enabling it to generate feasible trajectories under the current system. Additionally, by guiding the sampling process, the generated trajectories can satisfy additional constraints.
The technical details of the diffusion model employed are provided in the Appendix B. Here, we briefly outline the main idea. To satisfy the first condition, we control the length of the generated trajectory by adjusting the length of the initial noise during the denoising process. This is feasible due to the structural properties of the backbone network used in the diffusion model Janner et al. (2022). For the second condition, we treat it as an inpainting problem Janner et al. (2022), ensuring that the generated trajectory meets the requirement by replacing the start and end states of the trajectory with $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$, respectively, after each denoising step. To address the last condition, we decompose each invariance progress $\mathcal{I}(a,b,\mu_{i})$ into state constraints of the form $h_{\mu_{i}}(\mathbf{x}_{t})\geq 0$ for $t=a,a+1,\dots,b$, commonly referred to as “safety” constraints. We employ SafeDiffuser Xiao et al. (2023) to generate trajectories that comply with these safety constraints. SafeDiffuser integrates control barrier functions (CBFs) Nguyen and Sreenath (2016) to enforce finite-time diffusion invariance directly within the sampling process.
@@ 4 Case Study
Figure 3: Planned Trajectory (left) and Actual Execution Trajectory (right) in Case Study. The numbers next to start point and regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ indicate the completion times for each reachability progress assigned by progress allocation module.
@@
Table 1: STL Task Decomposition Results
To illustrate the workflow of our algorithm, we consider a sequential visit and region avoidance task in the Maze2D (Large) environment Fu et al. (2020).
In this scenario, the agent starts at the yellow point shown in Figure 3 and aims to complete the following STL task:
$\mathrm{F}_{[0,35]}(\mu_{1}\wedge(\mathrm{F}_{[35,45]}(\mu_{2}\wedge\mathrm{F}_{[10,30]}\mu_{3})))\wedge\mathrm{G}_{[0,110]}(\neg\mu_{4}\wedge\neg\mu_{5}),$
where the predicate $\mu_{i}$ represents “reach region $\mu_{i}$,” and $\neg\mu_{i}$ denotes “avoid region $\mu_{i}$”. Intuitively, this STL task requires the agent to sequentially visit circular regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ within specific time intervals while avoiding regions $\mu_{4}$ and $\mu_{5}$ throughout the entire episode. The environment layout and target regions are depicted in Figure 3. Our method leverages only an STL task-agnostic trajectory dataset to train the diffusion model, without prior knowledge of the map or system dynamics.
The decomposed progresses and time constraints are summarized in Table 1. The planning algorithm then assigns completion times to each reachability progress, as indicated by the numbers next to start point and regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ in Figure 3, resulting in a sequence of waypoints with corresponding times. The diffusion model is then used to sequentially generate trajectories between adjacent waypoints while ensuring all invariance progresses are satisfied. The final generated trajectory is shown in the left subfigure of Figure 3.
Finally, to evaluate the feasibility of the planned trajectory, we employ a simple model-free PD controller to track it. The resulting execution trajectory is shown in the right subfigure of Figure 3. Using the open-source library stlpy Kurtz and Lin (2022), we calculate the robustness values for both the planned and actual execution trajectories as 0.180 and 0.115, respectively. Since both values are positive, the trajectories satisfy the STL task requirements.

### v2

@@ 3.3 Progress Allocation
Prediction of Reachability Time. In Algorithm 2, model TimePredict is used to estimates the time step (trajectory length) needed to transition from the current state $\mathbf{x}$ to the new state $\mathbf{x}^{\prime}$. This model is trained on the same trajectory data, that will be used to train the Diffusion model. It assumes that the trajectory length between two states follows a Gaussian distribution. A simple multilayer perceptron (MLP) is used as the backbone of TimePredict model, and it is trained to predict the mean and variance of the trajectory length using a negative log-likelihood loss function. To account for tasks with avoidance requirements, which may require longer trajectories, a scaling factor $\gamma$ is applied to the predicted mean trajectory length. This factor allows control over the algorithm’s conservativeness by adjusting the predicted trajectory length as needed.
Soundness of the Progress Allocation Algorithm. We establish the soundness of the progress allocation algorithm in Lemma 2, with proof provided in Appendix G.
Lemma 2 (Soundness of Progress Allocation Algorithm).
Let $(\mathbf{x}_{0},\mathbb{P}^{\mathcal{R}},\mathbb{P}^{\mathcal{I}},\mathbb{T})$ be the input to Algorithm 1. Assume the depth-first allocation terminates with a waypoint sequence $\tilde{\mathbf{s}}=(\tilde{\mathbf{x}}_{0},t_{0})\,(\tilde{\mathbf{x}}_{1},t_{1})\,\dots\,(\tilde{\mathbf{x}}_{n},t_{n}),\quad 0=t_{0}\leq t_{1}\leq\dots\leq t_{n},$ and let $\mathbb{T}_{f}$ be the final set of time-variable constraints with feasible assignment space $\mathcal{F}_{f}$.
Then $\mathcal{F}_{f}\neq\varnothing$, and for any signal $\mathbf{s}_{0}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$ with $T\geq t_{n}$ and $\mathbf{x}_{t_{i}}=\tilde{\mathbf{x}}_{i}$ for $0\leq i\leq n$, there exists $\boldsymbol{\lambda}\in\mathcal{F}_{f}$ such that
$\left\{\begin{aligned} &\forall\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{R}}:\;\mathbf{s}_{0}\models\mathcal{R}(a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda}),\mu),\\ &\forall\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}},\;\forall t\in[a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]\cap\{t_{i}\}_{i=0}^{n}:\;\mathbf{x}_{t}\models\mu.\end{aligned}\right$.
In other words, by interpreting the returned waypoints as key states within the signal $\mathbf{s}_{0}$, the algorithm guarantees that, under some feasible assignment of time variables, all reachability progresses are satisfied and no invariance progress is violated at any selected waypoint.
@@ 3.4 Trajectory Generation
Lemma 2 ensures that the signal constructed from the waypoints returned by the Progress Allocation algorithm satisfies all reachability progresses and does not violate any invariance progress at the selected time steps. We now address the problem of completing the full system trajectory so that the entire signal satisfies all progress conditions. Suppose the Progress Allocation algorithm returns a waypoint sequence $\tilde{\mathbf{s}}=(\tilde{\mathbf{x}}_{0},t_{0})\,(\tilde{\mathbf{x}}_{1},t_{1})\,\dots\,(\tilde{\mathbf{x}}_{n},t_{n}),\quad 0=t_{0}\leq t_{1}\leq\dots\leq t_{n}$. Our goal is to generate a system trajectory $\boldsymbol{\tau}=\mathbf{x}_{0}\mathbf{a}_{0}\mathbf{x}_{1}\mathbf{a}_{1}\dots\mathbf{a}_{T-1}\mathbf{x}_{T}$ under system dynamics (1), whose corresponding signal $\mathbf{s}_{0}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$ satisfies:
$\left\{\begin{aligned} &T\geq t_{n},\\ &\mathbf{x}_{t_{i}}=\tilde{\mathbf{x}}_{i}\quad\text{for all }i\in\{0,\dots,n\},\\ &\forall\,\mathcal{I}\in\mathbb{P}^{\mathcal{I}},\;\mathbf{s}_{0}\models\mathcal{I}.\end{aligned}\right$. (6)
Note that the time intervals of all invariance progresses are now fixed by selecting a specific feasible assignment $\boldsymbol{\lambda}^{\star}\in\mathcal{F}_{f}$. A special case arises when certain invariance progresses have not yet terminated at time $t_{n}$. In this situation, we simply wait at $\tilde{\mathbf{x}}_{n}$ until all invariance progresses end, since according to Lemma 2, $\tilde{\mathbf{x}}_{n}$ satisfies all predicates corresponding to the active invariance progresses at that time.
To solve this conditional trajectory generation problem, we employ a diffusion model trained on task-agnostic trajectory data, which enables the generation of dynamically feasible trajectories. Moreover, by incorporating constraints into the sampling process, the generated trajectories can be aligned with additional task requirements. However, directly generating long trajectories is challenging, as they often deviate from the training distribution. To mitigate this, we adopt a trajectory stitching strategy, generating individual segments $\boldsymbol{\tau}_{i}$ between each pair of adjacent waypoints $(\mathbf{x}_{i},t_{i})$ and $(\mathbf{x}_{i+1},t_{i+1})$, and concatenating them to form the final trajectory $\boldsymbol{\tau}$. Specially, each segment must satisfy the following conditions: (i) the length is $t_{i+1}-t_{i}+1$, (ii) it starts at state $\mathbf{x}_{i}$ and ends at state $\mathbf{x}_{i+1}$, and (iii) it satisfies all invariance progresses active within the time interval $t_{i},t_{i+1}]$.
The diffusion model implementation is detailed in Appendix E. Briefly, the trajectory length is controlled via the length of initial noise, leveraging architectural properties of the backbone network [16]. Endpoint conditions are handled as an inpainting problem [16], by fixing the start and end states at every denoising step. To enforce invariance progresses, each $\mathcal{I}(a,b,\mu_{i})$ is translated into a set of pointwise state constraints $h_{\mu_{i}}(\mathbf{x}_{t})\geq 0$ for $t=a,\dots,b$, commonly referred to as safety constraints. These constraints can be handled using various existing constrained sampling methods for diffusion models [31, 29, 32, 33, 34, 35]. We employ SafeDiffuser [29], which integrates control barrier functions (CBFs) [36] to enforce these constraints within the sampling process, providing formal guarantees under mild conditions.
Soundness Analysis. We establish the soundness of the overall framework in the following theorem. The proof is provided in Appendix G.
Theorem 1 (Soundness of the Overall Planner).
Let $\varphi$ be an STL formula satisfying all assumptions in Sections 2.2 and 3.2. Assume that the Semantics-based Decomposition terminates with a finite set of progresses and constraints, the Progress Allocation algorithm returns a sequence of timed waypoints and a non-empty feasible assignment set, and the Trajectory Generation module produces a trajectory $\boldsymbol{\tau}$ whose signal $\mathbf{s}_{0}$ visits all timed waypoints and satisfies all active invariance progresses. Then the resulting signal satisfies the STL specification: $\mathbf{s}_{0}\vDash\varphi$.
3.5 Action Sequence and Control Protocol
Following mainstream diffusion-planner practice, we use the diffusion model solely to generate the state sequence; since action sequences are high-frequency and less smooth [17], we recover them via an learned inverse dynamics model [37] or with controllers that track the generated state sequence during execution. To guarantee strict adherence to the temporal windows imposed by STL tasks, we adopt a time-synchronous control protocol: each planning step is aligned with a fixed number $k$ of control steps, where $k$ is specified at task definition time to relate the planning horizon to the control frequency. At each planning step, the controller or inverse dynamics model therefore outputs exactly $k$ actions and executes $k$ control updates, ensuring that the temporal structure of the trajectory is preserved at runtime; unless otherwise stated, we set $k=1$, so each planning step corresponds to precisely one control step.
@@ 4 Case Study
Table 1: STL Task Decomposition Results.
@@
To illustrate the workflow of our algorithm, we consider a sequential visit and region avoidance task in the Maze2D (Large) environment [38]. In this scenario, the agent starts at the yellow point shown in Figure 2 and aims to complete the following STL task: $\mathrm{F}_{[0,35]}(\mu_{1}\wedge(\mathrm{F}_{[35,45]}(\mu_{2}\wedge\mathrm{F}_{[10,30]}\mu_{3})))\wedge\mathrm{G}_{[0,110]}(\neg\mu_{4}\wedge\neg\mu_{5})$, where the predicate $\mu_{i}$ represents “reach region $\mu_{i}$,” and $\neg\mu_{i}$ denotes “avoid region $\mu_{i}$”. Intuitively, this STL task requires the agent to sequentially visit circular regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ within specific time intervals while avoiding regions $\mu_{4}$ and $\mu_{5}$ throughout the entire episode.
The environment layout and target regions are depicted in Figure 2. Our method leverages only an STL task-agnostic trajectory dataset to train the diffusion model, without prior knowledge of the map or system dynamics. The decomposed progresses and time constraints are summarized in Table 1. The planning algorithm then assigns completion times to each reachability progress, as indicated by the numbers next to start point and regions $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ in Figure 2, resulting in a sequence of waypoints with corresponding times. The diffusion model is then used to sequentially generate trajectories between adjacent waypoints while ensuring all invariance progresses are satisfied. The final generated trajectory is shown in the left subfigure of Figure 2. To evaluate the feasibility of the planned trajectory, we employ a simple PD controller to track it as described in Section 3.5.
Figure 2: Planned Trajectory (left) and Actual Execution Trajectory (right) in Case Study.
The resulting execution trajectory is shown in the right subfigure of Figure 2. Using the open-source library stlpy [2], we calculate the robustness values for both the planned and actual execution trajectories as 0.180 and 0.115, respectively. Since both values are positive, the trajectories satisfy the STL task requirements.